# Rejection via Learning Density Ratios

### v1

@@ 1 Introduction
In classification, a confidence-based approach can be utilized, where a classifier is trained to output a “margin” which is used as a confidence score for rejection [7, 28, 57, 47, 44]. Given a threshold value, the model rejects whenever this confidence score is lower than the threshold value. A key component of these confidence-based approaches is that they rely on providing good class probability estimates $\Pr(\mathsf{Y}\mid\mathsf{X}=x)$ [48], i.e., being calibrated [57, 44]. Some approaches avoid explicit estimation of probabilities, however, these are usually restricted to binary classification [7, 28, 38].
@@
Another classification-rejection approach aim to simultaneously train a prediction and rejection model in tandem [16, 17, 44]. These approaches are theoretically driven by the construction of surrogate loss functions, but in the multiclass classification case many of these loss functions have been shown to not be suitable [44]. For the multiclass classification setting, one approach connects classification with rejection to cost-sensitive classification [14]. In practice, these classification-rejection approaches require models to be trained from scratch using their specific loss function and architecture — if there is an existing classifier for the dataset, it must be discarded. A recent approach proposes to learn a rejector on top of a pretrained classifier via surrogate loss functions [39].
@@
In this work, to learn rejectors we shift from a loss function perspective to a distributional perspective. Given a model and a corresponding loss function, we find a distribution where the model and loss performs “best” and compare it against the data input distribution to make rejection decisions (see Fig. 1). As such, the set of rejectors that we propose creates a rejection decision by considering the density ratio [55] between a “best” case (idealized) distribution and the data distribution, which can be thresholded by different values $\tau$ to provide different accuracy vs rejection percentage trade-offs. To learn these density ratios for rejection, we consider a risk minimization problem which is regularized by $\varphi$-divergences [1, 19]. We study a particular type of $\varphi$-divergences as our regularizer: the family of $\alpha$-divergences which generalizes the KL-divergence. To this end, one of our core contributions in this work is providing various methods for constructing and approximating idealized distributions, in particular those constructed by $\alpha$-divergences. The idealized distributions that we consider are connected to adversarial distributions examined in Distributionally Robust Optimization (DRO) [51] and the distributions learned in Generalized Variational Inference (GVI) [34]; and, as such, the closed formed solutions found for our idealized distribution have utility outside of the rejection setting. Furthermore, when utilizing the KL-divergence and Bayes optimal models, we recover the well known optimal rejection policies, i.e., Chow’s rule [15, 58]. Our rejectors are then examined empirically by examining the classification learning setting with calibrated classifiers.
@@
  • We present a set of simplifying assumptions such that our framework can be utilized in practice and verify this empirically.
$x$Data Distribution $\mathrm{P}$Idealized Distribution $\mathrm{Q}$Reject if $\rho=\frac{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{d}\mathrm{Q}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{d}\mathrm{P}}(x)\leq\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\tau$
Figure 1: We first learn an idealized distribution $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Q}$ which minimizes the loss function of a model. We then compare the $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Q}$ with the original data distribution $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{P}$ via calculating the density ratio $\rho={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{d}\mathrm{Q}}/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{d}\mathrm{P}}$. This is then used to define a rejection criteria via threshold value $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\tau$.
@@
Let $\mathcal{X}$ be a domain of inputs and $\mathcal{Y}$ be output targets. We primarily consider the case where $\mathcal{Y}$ is a finite set of $N$ labels $[N]$, where $[N]\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\{0,\ldots,N-1\}$. We will also consider output domains $\mathcal{Y}^{\prime}$ which is not necessarily the same as the output data $\mathcal{Y}$, e.g., class probabilities estimates in the classification setting. We further denote the underlying (usually unknown) input distribution as $\mathrm{P}_{\rm x,\rm y}\in\mathcal{D}(\mathcal{X}\times\mathcal{Y})$. The marginals of a joint distribution are denoted by subscript, e.g., $\mathrm{P}_{\rm y}\in\mathcal{D}(\mathcal{Y})$ for the marginal on the label space. Similarly, we denote conditional distributions, e.g., $\mathrm{P}_{\rm x\mid\rm y}$. For marginals on $\mathcal{D}(\mathcal{X})$, we make the subscripting of $\rm x$ implicit with $\mathrm{P}=\mathrm{P}_{\rm x}$. The empirical distributions of $\mathrm{P}$ are denoted by $\hat{\mathrm{P}}_{N}$, with $N$ denoting the number of samples used to generate it. Furthermore, denote the Iverson bracket $\llbracket p\rrbracket=1$ if the predicate $p$ is true and $\llbracket p\rrbracket=0$ otherwise [35]. The maximum $[z]_{+}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\max\{0,z\}$ is shorthanded.
2 Preliminaries
@@
One way of viewing learning with rejection is to augment the risk minimization framework by learning an addition rejection function $r\colon\mathcal{X}\rightarrow\{0,1\}$. This can be formally defined using a regularization term $c\in\mathbb{R}_{\geq 0}$ which controls the rate of rejection:
@@
where $\mathcal{R}$ denotes a hypothesis set of rejection functions.
@@
Once minimized, a combined model with rejection can be defined by,
$f(x)=\begin{cases}\circledR&\textrm{if }r(x)=1\\ h(x)&\textrm{if }r(x)=0,\end{cases}$ (3)
[…]
@@
Let us consider the binary CPE setting, where $\mathcal{Y}=\{0,1\}$, $\mathcal{Y}^{\prime}=[0,1]$, and $\ell$ be any proper [^1] Properness ensures that the true class probability is the minimizer of the loss in expectation. loss function [48] (e.g., log loss). Then w.r.t. Eq. 2, the optimal classifier is given by $h^{\star}(x)=\mathrm{P}(\mathsf{Y}=1\mid\mathsf{X}=x)$ and the optimal rejector is given by $r^{\star}(x)=\llbracket\mathbb{E}_{\mathsf{Y}\sim{h^{\star}(\mathsf{X}=x)}}[\ell(\mathsf{Y},h^{\star}(x))]\geq c\rrbracket$.
The theorem can easily be generalized to non-binary cases. We note that Theorem 2.1 is a generalization of the well known Chow’s rule of classification with rejection [15, 14]. Indeed, taking $\ell$ to be the zero-one-loss function, we get $r^{\star}(x)=\llbracket|2\cdot\mathrm{P}(\mathsf{Y}=1\mid\mathsf{X}=x)-1|\geq 1-c\rrbracket$. One can further clarify this by noticing that $\Pr[r(\mathsf{X})=1]=\mathbb{E}_{\mathrm{P}}[r(\mathsf{X})]$. In the general proper loss case, the optimal rejector is thresholding a generalized entropy function (known as the conditional Bayes risk [48]). This would also correspond to thresholding the class probabilities $\mathrm{P}(\mathsf{Y}=y\mid\mathsf{X}=x)$, but with different thresholding values per class $y\in\mathcal{Y}$ (unless $\ell$ is a symmetric loss function).
@@
Generalized Variational Inference (GVI) [34] provides a framework for a generalized set of entropy regularized risk minimization problem. In particular, GVI generalizes Bayes’ rule by interpreting the Bayesian posterior as the solution to a minimization problem [59]. The Bayes’ rule minimization is given by,
@@
A related piece of literature, is Distributionally Robust Optimization (DRO) [51], where the goal is to find a distribution that maximizes (or minimizes) the expectation of a function from a prescribed uncertainty set. Popular candidates for these uncertainty sets include all distributions that are a certain radius away from a fixed distribution by some divergence. $\varphi$-divergences have been used to define such uncertainty set [8, 22, 36]. Given radius $\varepsilon>0$, define $B^{\varphi}_{\varepsilon}({\mathrm{P}})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\left\{{\mathrm{Q}}\in\mathcal{D}(\mathcal{X}\times\mathcal{Y}):D_{\varphi}({\mathrm{P}}\mathrel{\|}{\mathrm{Q}})<\varepsilon\right\}$. Given a point-wise loss function $\ell\colon\mathcal{Y}\times\mathcal{Y}^{\prime}\rightarrow\mathbb{R}$, DRO alters risk minimization, as per Eq. 1, to solve the following optimization problem:
@@
The maximization over $\bar{\mathrm{Q}}$ is typically over the target space $\mathcal{Y}$. That is, $\bar{\mathrm{Q}}(x,y)=\mathrm{Q}(y)\cdot\mathrm{P}(x\mid y)$ and the maximization is adversarial over the marginal label distribution [60]. We note that by converting the $\varepsilon$-ball constraint into a Lagrange multiplier, the inner optimization over $\mathrm{Q}$ in Eq. 7 becomes a optimization problem which mirrors Eq. 5. The connection between GVI and adversarial robustness has been previously noted [31].
@@ 3 Rejection via Idealized Distributions
where $a(x)$ are Lagrange multipliers to ensure non-negativity $\rho^{\varphi}(\cdot)\geq 0$; and $b$ is a Lagrange multiplier to ensure the constraint $\mathbb{E}_{\mathrm{P}}\left[\rho^{\varphi}_{\lambda}(\mathsf{X})\right]=1$ is satisfied. Furthermore, the optimal idealized rejection distribution is given by: $\mathrm{Q}^{\varphi}(x)=\mathrm{P}(x)\cdot\rho^{\varphi}_{\lambda}(x)$.
@@ 4 Learning Idealized Distributions
KL Divergence
Let us first consider the KL-divergence [2] for constructing density ratio rejectors.
@@
Let $\varphi(z)=z\log z-z+1$ and $\lambda>0$. The optimal density ratio of Definition 3.1 is,
$\rho_{\lambda}^{\textsc{KL}}(x)$ $=\frac{1}{Z}\cdot\exp\left(\frac{-L^{\prime}(x)}{\lambda}\right),\quad\quad\textrm{where }Z=\mathbb{E}_{\mathrm{P}}\left[\exp\left(\frac{-L^{\prime}(x)}{\lambda}\right)\right].$ (13)
One will notice that $\rho^{\textsc{KL}}_{\lambda}$ corresponds to an exponential tilt [24] of $\mathrm{P}$ to yield a Gibbs distribution $\mathrm{Q}^{\textsc{KL}}$. The KL density ratio rejectors are significant in a few ways. First, from a closed-form solution perspective, due to the properties of ‘$\log$’ and complementary slackness, $\alpha(y)=0$. Secondly, due to the properties of $\exp$, the normalization term $b$ has a closed form solution given by the typical log-normalizer term of exponential families.

### v2

@@ 1 Introduction
(†) Work done while interning at Amazon.(*) Work done while employed at Amazon.
@@
In classification, a confidence-based approach can be utilized, where a classifier is trained to output a “margin” which is used as a confidence score for rejection [7, 31, 71, 59, 53]. The model rejects whenever this confidence score is lower than an assigned threshold value. A key aspect of these approaches is that they rely on good probability estimates $\Pr(\mathsf{Y}\mid\mathsf{X}=x)$ [60], i.e., being calibrated [71, 53]. While some approaches avoid explicit probability estimation, these are typically restricted to binary classification [7, 31, 47]. Empirically, approaches utilizing confidence scores have shown to outperform other methods, even with simple plugin estimates for probabilities [26, 39, 58].
@@
Another classifier-rejection approach aims to simultaneously train a prediction and rejection model in tandem [16, 17, 53]. These approaches are theoretically driven by the construction of surrogate loss functions, but in the multiclass classification case many of these loss functions have been shown to not be suitable [53]. For the multiclass classification setting, one approach connects classification with rejection to cost-sensitive classification [14]. In practice, these classification-rejection approaches require models to be trained from scratch using their specific loss function and architecture — if there is an existing classifier for the dataset, it must be discarded. A recent approach proposes to learn a post-hoc rejector on top of a pretrained classifier via surrogate loss functions [48].
@@
In this work, to learn rejectors we shift from a loss function perspective to a distributional perspective. Given a model and a corresponding loss function, we find a distribution where the model and loss performs “best” and compare it against the data input distribution to make rejection decisions (see Fig. 1 and Algorithm 1 with equations boxed). As such, the set of rejectors that we propose creates a rejection decision by considering the density ratio [67] between a “best” case (idealized) distribution and the data distribution, which can be thresholded by different values $\tau$ to provide different accuracy vs rejection percentage trade-offs. To learn these density ratios for rejection, we consider a risk minimization problem which is regularized by $\varphi$-divergences [1, 19]. We study a particular type of $\varphi$-divergences as our regularizer: the family of $\alpha$-divergences which generalizes the KL-divergence. To this end, one of our core contributions in this work is providing various methods for constructing and approximating idealized distributions, in particular those constructed by $\alpha$-divergences.
The idealized distributions that we consider are connected to adversarial distributions examined in Distributionally Robust Optimization (DRO) [63] and the distributions learned in Generalized Variational Inference (GVI) [42]; and, as such, the closed formed solutions found for our idealized distribution have utility outside of rejection. Furthermore, when utilizing the KL-divergence and Bayes optimal models, we recover the well known optimal rejection policies, i.e., Chow’s rule [15, 72]. Our rejectors are then examined empirically on 6 different datasets and under label noise corruption.
@@
  • We present a set of simplifying assumptions such that our framework can be utilized in practice for post-hoc rejection and verify this empirically.
$x$Data Dist. $\mathrm{P}$Idealized Dist. $\mathrm{Q}$Reject if $\rho(x)=\frac{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{d}\mathrm{Q}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{d}\mathrm{P}}(x)\leq\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\tau$
Figure 1: An idealized distribution $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Q}$ is learned to minimizes the loss of a model. We then compare $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Q}$ with the original data distribution $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{P}$ via a density ratio $\rho={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{d}\mathrm{Q}}/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{d}\mathrm{P}}$. A rejection criteria is defined via threshold value $\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\tau$.
Algorithm 1 Density-Ratio Rejection
0:  Model $h$; Divergence $\alpha\geq 1$; Reg. $\lambda>0$; Threshold $\tau\in[0,1]$.
1:  Calibrate $h$ if needed.
2:  Calculate density ratio $\rho_{\lambda}^{\alpha}$ (either Corollary 4.1 or 4.6).
3:  Normalize $\rho_{\lambda}^{\alpha}$ with Monte-Carlo or Eq. 17.
4:  Calculate rejector $r_{\tau}$ via Def. 3.2.
4:  $r_{\tau}$
2 Preliminaries
@@
Let $\mathcal{X}$ be a domain of inputs and $\mathcal{Y}$ be output targets. We primarily consider the case where $\mathcal{Y}$ is a finite set of $N$ labels $[N]$, where $[N]\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\{0,\ldots,N-1\}$. We also consider output domains $\mathcal{Y}^{\prime}$ which is not necessarily the same as the output data $\mathcal{Y}$, e.g., class probabilities estimates in classification. Denote the underlying (usually unknown) input distribution as $\mathrm{P}_{\rm x,\rm y}\in\mathcal{D}(\mathcal{X}\times\mathcal{Y})$. The marginals of a joint distribution are denoted by subscript, e.g., $\mathrm{P}_{\rm y}\in\mathcal{D}(\mathcal{Y})$ for the marginal on the label space. We denote conditional distributions, e.g., $\mathrm{P}_{\rm x\mid\rm y}$. For marginals on $\mathcal{D}(\mathcal{X})$, the subscripting of $\rm x$ is implicit with $\mathrm{P}=\mathrm{P}_{\rm x}$. The empirical distributions of $\mathrm{P}$ are denoted by $\hat{\mathrm{P}}_{N}$, with $N$ denoting the number of samples used to generate it. Denote the Iverson bracket $\llbracket p\rrbracket=1$ if the predicate $p$ is true and $\llbracket p\rrbracket=0$ otherwise [43]. The maximum $[z]_{+}\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\max\{0,z\}$ is shorthanded.
@@
One way of viewing learning with rejection is to augment the risk minimization framework by learning an additional rejection function $r\colon\mathcal{X}\rightarrow\{0,1\}$. This can be formally defined using a regularization term $c\in\mathbb{R}_{\geq 0}$ which controls the rate of rejection:
@@
[…]
$f(x)=\begin{cases}\circledR&\textrm{if }r(x)=1\\ h(x)&\textrm{if }r(x)=0.\end{cases}$ (3)
@@
Let us consider the binary CPE setting, where $\mathcal{Y}=\{0,1\}$, $\mathcal{Y}^{\prime}=[0,1]$, and $\ell$ be any proper [^1] Properness ensures that the true class probability is the minimizer of the loss in expectation. loss function [60] (e.g., log loss). Then w.r.t. Eq. 2, the optimal classifier is given by $h^{\star}(x)=\mathrm{P}[\mathsf{Y}=1\mid\mathsf{X}=x]$ and the optimal rejector is given by $r^{\star}(x)=\llbracket\mathbb{E}_{\mathsf{Y}\sim{h^{\star}(\mathsf{X}=x)}}[\ell(\mathsf{Y},h^{\star}(x))]\geq c\rrbracket$.
The theorem can easily be generalized to non-binary cases. We note that Theorem 2.1 is a generalization of the well known Chow’s rule of classification with rejection [15, 14]. Indeed, taking $\ell$ to be the zero-one-loss function, we get $r^{\star}(x)=\llbracket|2\cdot\mathrm{P}[\mathsf{Y}=1\mid\mathsf{X}=x]-1|\geq 1-c\rrbracket$. One can further clarify this by noticing that $\Pr[r(\mathsf{X})=1]=\mathbb{E}_{\mathrm{P}}[r(\mathsf{X})]$. In the general proper loss case, the optimal rejector is thresholding a generalized entropy function (known as the conditional Bayes risk [60]). This would correspond to thresholding the class probabilities $\mathrm{P}[\mathsf{Y}=y\mid\mathsf{X}=x]$ (via the Bayes posterior $h^{\star}$), but with different thresholding values per class $y\in\mathcal{Y}$ (unless $\ell$ is a symmetric loss function).
Although the objective of Eq. 2 follows the cost-based model of rejection [15], other models of rejection exist in the literature. Alternatively, the bounded improvement model of rejection [54, 29] maximizes coverage (non-rejection rate) whilst maintaining a constraint on the performance of the rejection measured by the selective risk (the first term of Eq. 2 inversely weighted by the coverage). Optimality conditions of the bounded improvement have been explored in [27]. In the bounded abstention model, the constraint and objective is switched – the selective risk is minimized with a constraint on minimum coverage [54]. Our objective function Eq. 2 (and the cost-based model) can be interpreted as a change in the type of risk considered and a change in the hard coverage constraint to a soft constraint w.r.t. bounded abstention. The optimal strategies of these approaches are explored in Franc et al. [28]. Alongside Section 1, further details about rejection can be found in Hendrickx et al. [34].
@@
Generalized Variational Inference (GVI) [42] provides a framework for a generalized set of entropy regularized risk minimization problems. In particular, GVI generalizes Bayes’ rule by interpreting the Bayesian posterior as the solution to a minimization problem [73]. The Bayes’ rule minimization is given by,
@@
A related piece of literature is Distributionally Robust Optimization (DRO) [63], where the goal is to find a distribution that maximizes (or minimizes) the expectation of a function from a prescribed uncertainty set. Popular candidates for these uncertainty sets include all distributions that are a certain radius away from a fixed distribution by some divergence. $\varphi$-divergences have been used to define such uncertainty set [8, 22, 45]. Given radius $\varepsilon>0$, define $B^{\varphi}_{\varepsilon}({\mathrm{P}})\stackrel{{\scriptstyle\mathrm{.}}}{{=}}\left\{{\mathrm{Q}}\in\mathcal{D}(\mathcal{X}\times\mathcal{Y}):D_{\varphi}({\mathrm{P}}\mathrel{\|}{\mathrm{Q}})<\varepsilon\right\}$. Given a point-wise loss function $\ell\colon\mathcal{Y}\times\mathcal{Y}^{\prime}\rightarrow\mathbb{R}$, DRO alters risk minimization, as per Eq. 1, to solve the following optimization problem:
@@
The max over $\bar{\mathrm{Q}}$ is typically over the target space $\mathcal{Y}$. That is, $\bar{\mathrm{Q}}(x,y)=\mathrm{Q}(y)\cdot\mathrm{P}(x\mid y)$ and the max is adversarial over the marginal label distribution [74]. Note that converting the $\varepsilon$-ball constraint into a Lagrange multiplier, the inner optimization over $\mathrm{Q}$ in Eq. 7 mirrors Eq. 5. The connection between GVI and adversarial robustness has been previously noted [37].
@@ 3 Rejection via Idealized Distributions
where $a(x)$ are Lagrange multipliers to ensure non-negativity $\rho_{\lambda}^{\varphi}(\cdot)\geq 0$; and $b$ is a Lagrange multiplier to ensure the constraint $\mathbb{E}_{\mathrm{P}}\left[\rho^{\varphi}_{\lambda}(\mathsf{X})\right]=1$ is satisfied. Furthermore, the optimal idealized rejection distribution is given by: $\mathrm{Q}^{\varphi}(x)=\mathrm{P}(x)\cdot\rho^{\varphi}_{\lambda}(x)$.
Taking $h$ as the Bayes posterior $h^{\star}$, $L^{\prime}$ becomes a function of the ground truth posterior $\Pr(\mathsf{Y}\mid\mathsf{X}=x)$. Hence taking the output $h$ as a neural network plugin estimate of the underlying true posterior $\Pr(\mathsf{Y}\mid\mathsf{X}=x)$ (see Section 4.3) yields an approach similar to softmax response, i.e., rejection based on the output of $h$ when it outputs softmax probabilities [29]. Theorem 3.3 presents a general approach to generating rejectors from these plugin estimates, as a function of $\ell$ and $\varphi$.
@@ 4 Learning Idealized Distributions
4.1 KL-Divergence
Let us first consider the KL-divergence [2] for constructing density ratio rejectors via Theorem 3.3.
@@ 4.1 KL-Divergence
Let $\varphi(z)=z\log z-z+1$ and $\lambda>0$. The optimal density ratio of Definition 3.1 is, $\rho_{\lambda}^{\textsc{KL}}(x)$ $=\frac{1}{Z}\cdot\exp\left(\frac{-L^{\prime}(x)}{\lambda}\right),\quad\quad\textrm{where }Z=\mathbb{E}_{\mathrm{P}}\left[\exp\left(\frac{-L^{\prime}(x)}{\lambda}\right)\right].$ (13)
One will notice that $\rho^{\textsc{KL}}_{\lambda}$ corresponds to an exponential tilt [24] of $\mathrm{P}$ to yield a Gibbs distribution $\mathrm{Q}^{\textsc{KL}}$. The KL density ratio rejectors are significant in a few ways. First, we obtain a closed-form solution due to the properties of ‘$\log$’ and complementary slackness, i.e., $a(\cdot)=0$. Secondly, due to the properties of $\exp$, the normalization term $b$ has a closed form solution given by the typical log-normalizer term of exponential families.